# The Vizier Gaussian Process Bandit Algorithm

### v1

@@ 3.1.2 Output Preprocessing
  • Linear Scaling: Let $\xi(I):=\left(\sum\left\{(y_{i}-y_{\operatorname{median}})^{2}:i\in I\right\}\right)^{1/2}$. Divide by $\xi(\left\{i:y_{i}\geq y_{\operatorname{median}}\right\})$ if non–zero, else by $\xi(\left\{i:1\leq i\leq t\right\})$ if non–zero. Finally, shift the data so the median is zero.
@@
  • Log Warping: Increases modeling resolution for good values. We first normalize all objectives as $\widehat{y}\leftarrow\frac{y_{max}-y}{y_{max}-y_{min}}$ and then define the warped values as $\widehat{y}\leftarrow 0.5-\log\big{(}1+\widehat{y}_{i}(s-1)\big{)}/{\log(s)}$ where $s$ is a free parameter set to $1.5$. This gently stretches the intervals between better y-values and compresses the intervals between worse values.
  • Infeasibility Warping: We replace all infeasible objectives with $y_{min}-(0.5(y_{max}-y_{min}))$. This maps infeasible objectives to a reasonably scaled unpromising value to prevent the optimizer from re-exploring infeasible regions.
@@ 3.4 Acquisition Function and Trust Regions
The penalty value $-10^{12}$ is guaranteed to be lower than $\text{UCB}(x)$ due to our preprocessing steps and zero mean function. The additional distance term in the penalty is needed to provide gradient directions for acquistion optimizers to move into the trust region.
@@ 3.5.1 Firefly Algorithm
In order to use the Firefly algorithm an acquisition optimizer, we modified it with three crucial improvements:
@@ 4 Experiments: Ray Tune Comparisons
We compare the Vizier algorithm with well-established baselines published on Ray Tune (Liaw et al., 2018): Ax (Balandat et al., 2020), BayesianOptimization (Nogueira, 2014), HEBO (Cowen-Rivers et al., 2022), HyperOpt (Bergstra et al., 2015), Optuna (Akiba et al., 2019), and Scikit-Optimize (Head et al., 2018). These baselines are all variants of Bayesian optimization, but mostly differ in their acquisition functions and optimizers. A summary of these differences are in Appendix C.2.
@@ C.2 Algorithm Summaries
Ax: Depending on search space, the algorithm is selected by choose_generation_strategy [^3] https://github.com/facebook/Ax/blob/main/ax/modelbridge/dispatch_utils.py. The number of continuous vs. categorical parameters, affects the choice of GP model (Eriksson and Jankowiak, 2021) and kernel. The default acquisition is an expected improvement (EI)-based qNoisyExpectedImprovement. Note we do not provide objective noise variance in advance. For acquisition optimization, L-BFGS-B is used for continuous spaces while a sequential greedy algorithm is used for mixed and categorical spaces.

### v3

@@ 3.1.2 Output Preprocessing
  • Linear Scaling: Let $\xi(I):=\sqrt{\frac{1}{|I|}\sum_{i\in I}(y_{i}-y_{\operatorname{median}})^{2}}$. Let our deviation be $\xi(\left\{i:y_{i}\geq y_{\operatorname{median}}\right\})$ if non–zero, else by $\xi(\left\{i:1\leq i\leq t\right\})$ if non–zero. Shift the y-values so the median is zero, and then divide by the deviation.
@@
  • Log Warping: Increases modeling resolution for good values. We first apply the normalization $\widehat{y}\leftarrow\frac{y_{max}-y}{y_{max}-y_{min}}$ and then the warping $\widehat{y}\leftarrow 0.5-\log(1+(s-1)\cdot\widehat{y})/{\log(s)}$ where $s$ is a free parameter set to $1.5$. This gently stretches the intervals between better y-values and compresses the intervals between worse values.
  • Infeasibility Warping: We replace all infeasible objectives with $y_{min}-0.5(y_{max}-y_{min})$. This maps infeasible objectives to a reasonably scaled unpromising value to prevent the optimizer from re-exploring infeasible regions.
@@ 3.4 Acquisition Function and Trust Regions
The penalty value $-10^{12}$ is guaranteed to be lower than $\text{UCB}(x)$ due to our preprocessing steps and zero mean function. The additional distance term in the penalty is needed to provide ascent directions for acquistion optimizers to move into the trust region.
@@ 3.5.1 Firefly Algorithm
In order to use the Firefly algorithm as an acquisition optimizer, we modified it with three crucial improvements:
@@ 4 Experiments: Ray Tune Comparisons
We compare the Vizier algorithm with well-established baselines published on Ray Tune (Liaw et al., 2018): Ax/BoTorch (Balandat et al., 2020), BayesianOptimization (Nogueira, 2014), HEBO (Cowen-Rivers et al., 2022), HyperOpt (Bergstra et al., 2015), Optuna (Akiba et al., 2019), and Scikit-Optimize (Head et al., 2018). These baselines are all variants of Bayesian optimization, but mostly differ in their acquisition functions and optimizers. A summary of these differences are in Appendix C.2.
@@ C.2 Algorithm Summaries
Ax/BoTorch: Depending on search space, choose_generation_strategy [^3] https://github.com/facebook/Ax/blob/main/ax/modelbridge/dispatch_utils.py chooses the algorithm. The number of continuous vs. categorical parameters, affects the choice of GP model (Eriksson and Jankowiak, 2021) and kernel. The default acquisition is an expected improvement (EI)-based qNoisyExpectedImprovement. Note we do not provide objective noise variance in advance. For acquisition optimization, L-BFGS-B is used for continuous spaces while a sequential greedy algorithm is used for mixed and categorical spaces.